OpenStudy (anonymous):

A box of rotini noodles contains 10 servings. If you eat 7/8 of the recommended serving as a meal, how many complete meals can you get from one box?

OpenStudy (anonymous):

Well did you get an answer yet?

OpenStudy (anonymous):

NOPE @blahblah_who_cares

OpenStudy (anonymous):

Do you think you might have an idea on how to get it?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

Ok. the equation your asking is similar to this one. If i have 10 apples and I eat 1, how many times can I eat the ten apples. So what method of mathmatics do you use to solve this equation?

OpenStudy (ivancsc1996):

Bueno, eso significa que el contenido total (T) sobre el numero de servidas (n) es el contenido de cada servida(x). \[\frac{ T }{ n }=x \rightarrow \frac{ 7 }{ 8 }x \rightarrow ?n \rightarrow n=\frac{ 8T }{ 7x }\]

OpenStudy (ivancsc1996):

?

OpenStudy (ivancsc1996):

Ohhhhhhhh

OpenStudy (anonymous):

NO LO ENTIENDO NO SON LOS RESULTADOS QUE ME DAN

OpenStudy (ivancsc1996):

No, por el lado de blahblah es mejor, olvida lo que escribí arriba. Dame un segundo.

OpenStudy (anonymous):

A11 c.8.75 b.8 d.11.43

OpenStudy (anonymous):

what are u teaching him? Before you get confused, all you need to do is divide the ten servings by 7/8 that you had.

OpenStudy (ivancsc1996):

11.43 es la respuesta

OpenStudy (anonymous):

MUCHAS GRACIAS :) @ivancsc1996

OpenStudy (ivancsc1996):

I did it with conversion factor.\[10x=T \rightarrow 1y=\frac{ 7 }{ 8 }x \rightarrow 10x \frac{ 1y }{ \frac{ 7 }{ 8 }x }=11.43y=\] were 1y is one serving of the person, 1x is one recommended serving and T is the total amount

OpenStudy (anonymous):

@ivancsc1996 you need to teach him not give him answers. lol i thought u only knew spanish so i was slowly typing spanish to say that.

OpenStudy (ivancsc1996):

Jajajaja you are right.

OpenStudy (anonymous):

@ivancsc1996 dont worry. I used to give out answers until i realized that many people take advanatage to let others do their work. By teaching you can eliminate that and have people actually learn so they can help others.

OpenStudy (anonymous):

its not D Guys! @ivancsc1996

OpenStudy (ivancsc1996):

Damn then I don't know.

OpenStudy (anonymous):

hahaha maybe 11???